OpenStudy (anonymous):

Classify the following function: f(x) = 25x^2 - 1 Quadratic Linear Exponential None of the above

OpenStudy (anonymous):

What do you think it could be?

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

I'd say C

OpenStudy (princeharryyy):

quad

OpenStudy (anonymous):

linear equations are usually in this form: y = mx + b quadratics: y = ax^2 + bx + c exponential: y = a^x

OpenStudy (anonymous):

@princeharryyy don't just give out the answer..

OpenStudy (princeharryyy):

@Keridavis

OpenStudy (anonymous):

Thank you Evil.

OpenStudy (princeharryyy):

ohk @eviI u already gave the explanation but there are many others

OpenStudy (anonymous):

So it would be A?

OpenStudy (princeharryyy):

yup